OpenStudy (lgbasallote):

\[\huge \mathcal L \lbrace e^{-at} - e^{-bt} \rbrace\]sounds like a trick question....

OpenStudy (dominusscholae):

Hint: Above expression equals \[L(e^{-at}) - L(e^{-bt}) \]

OpenStudy (lgbasallote):

that's it?

OpenStudy (lgbasallote):

how do i get hte limits? the s> thingies?

OpenStudy (dominusscholae):

Upper bound = infinity Lower Bound: 0. Nothing about the problem says anything about special conditions for the bounds, so assume that you use the normal ones.

OpenStudy (lgbasallote):

no not that

OpenStudy (lgbasallote):

the s > thingies

OpenStudy (dominusscholae):

....Clarify?

OpenStudy (lgbasallote):

i dont know what it's called...like \[\mathcal L \lbrace \cos kt \rbrace = \frac{s}{s^2 + k^2}; \; \text{for} \; s>0\] that s > thingy

OpenStudy (dominusscholae):

Oh....wait then its s > 0 like normal. For the most part, its always s > 0 for just one function you're trying to transform. It will be different if, let's say, you have a piecewise function and there are different bounds for each equation.

OpenStudy (lgbasallote):

no..it's not s>0 :/

OpenStudy (dominusscholae):

Wait, ok now I get it. First, do the laplace transform. Then find which region the transform is UNDEFINED or DOESN"T EXIST.

OpenStudy (dominusscholae):

The transform simplifies to \[1/(s+a) - 1/(s+b) = (b-a)/{(s+a)(s+b)}\]

OpenStudy (dominusscholae):

So its undefined when s = a or s = b. So the answer should be: All real numbers s expect a and b.

OpenStudy (lgbasallote):

my book says s>max(-a, -b) idk what that means

OpenStudy (dominusscholae):

Oops sorry I mean "except -a and -b". This basically means either -a or -b depending on which is the largest.

OpenStudy (lgbasallote):

what do you call that s> thingy?

OpenStudy (dominusscholae):

"All values of s greater than"

OpenStudy (lgbasallote):

no i meant i was calling it "limits thingy" earlier but you didnt understand it so what is it actually called?

OpenStudy (dominusscholae):

Umm I don't know. I don't think it's a formal word for this.

OpenStudy (lgbasallote):

oh well thanks

OpenStudy (anonymous):

\(\mathsf{s}\) is here frequency domain I guess...

OpenStudy (anonymous):

*frequency only..